# Charged Nariai black holes on the dark bubble

### v1

@@ 1 Introduction
According to the dark bubble model, the universe is riding an expanding bubble in AdS5, with the radial direction along the throat Banerjee:2018qey . The bubble is formed through a phase transition, where a bubble of true vacuum nucleates. Its wall consists of a brane with a tension that is slightly less than the critical value, with the difference setting the value of the positive cosmological constant. The model has been studied and developed in a series of papers, see refs. Banerjee:2018qey ; Banerjee:2019fzz ; Banerjee:2020wix ; Banerjee:2020wov ; Banerjee:2021qei ; Danielsson:2022fhd ; Danielsson:2022lsl ; Danielsson:2023alz ; Basile:2023tvh ; Banerjee:2023uto , with the earlier results reviewed in refs. Banerjee:2021yrb ; Banerjee:2022ree and related studies in Berglund:2022qsb ; Bandos:2023yyo ; Koga:2019yzj ; Koga:2020jok ; Koga:2022opd ; Basile:2020mpt ; Basile:2021vxh . In Danielsson:2022lsl ; Danielsson:2023alz the dark bubble model was embedded into a specific string theory setting, where the unique hierarchy of scales of the model was explored. In standard compactifications, as well as in RS-brane worlds, the 5d Planck length is always larger than the 4d one, provided that the size of the extra dimension, and the AdS length scale, are larger than the higher dimensional Planck length. The dark bubble, which has a brane with an inside and an outside, rather than the two insides of the RS-brane, is fundamentally different. What matters is the difference in AdS scale between the inside and the outside, which leads to a smaller 5d Planck length than the 4d one. This leads to several unique phenomenological features - including no need for scale separation in the higher dimensions.
@@
In this paper, we consider another case that further tests the dark bubble model: a charged Nariai black hole as large as the cosmological horizon. Such black holes have been studied in the context of the Swampland paradigm, Montero:2019ekk , where it has been argued that their existence leads to new constraints on particle masses beyond those of the weak gravity conjecture. In this paper, we show that the dark bubble scenario can successfully incorporate a Nariai black hole. The analysis provides yet another explicit example of how the dark bubble reproduces 4d Einstein gravity without localising it on the brane, a characteristic of the RS model. Instead, the bulk plays an important role in the construction, and the failure to localise yields well-controlled corrections that become important only at stringy densities, Basile:2023tvh .
@@ 2 The dark bubble scenario
and $e^{\alpha}_{\;b}$ are the tangent vectors to the brane for every $b$. The “--” and “$+$” correspond to the contributions from the in-in- and outside spacetimes respectively. The 4d Newton’s constant is related to the 5d one through
@@
An important property, peculiar to the dark bubble, is that $S_{ab}$ seemingly contributes with the wrong sign to the 4d energy density. This is the minus sign in front of $S_{ab}$ in (2). In the case of a pure brane, we simply have $S_{ab}=-\sigma h_{ab}$, which is why a larger value of the brane density $\sigma$ reduces the effective cosmological constant given by $\Lambda_{4}=\sigma_{c}-\sigma$ [^2] Note that we are using -$+$ signature so that $h_{tt}<0$ has $S_{tt}>0$ (and $S^{t}_{t}<0$) for a positive energy density. For the dark bubble to nucleate, its tension needs to be smaller than the critical value $\sigma_{c}$, guaranteeing a positive cosmological constant.
To get the full energy-momentum tensor $T_{ab}$ in the presence of matter on the brane, one must take into account its backreaction on the bulk, captured by $\mathcal{J}_{ab}$. This gives rise to an effective total energy density in 4d that is equal to the physical one. It is this phenomenon that has been exhibited in earlier work mentioned in the introduction. The choice of the bulk solution is part of setting up the appropriate initial conditions for the physical situation of interest. It is a non-trivial result of ref. Basile:2023tvh that this can be done for e.g. electromagnetic waves such that the subsequent time evolution is that of general relativity coupled to Maxwell theory. A crucial aspect is how the electromagnetic field on the brane sources a string $B$-field in the bulk. The backreaction of this field on the bulk metric is what allows the tensor $\mathcal{J}_{ab}$ to give rise to the desired energy-momentum tensor $T_{ab}$.
@@ 3 Reissner-Nordström-de Sitter black holes and the Nariai limit
In this section, we briefly review Reissner-Nordström-de Sitter (RN-dS) black holes and the Nariai limit in four dimensions. These have been of attention recently and led to the Festina Lente bound in Montero:2019ekk . These black holes are solutions of Einstein-Maxwell theory given by the familiar action
@@
The constraint (13) and the fact that $\hat{r}_{-}\leq\hat{r}_{+}\leq\hat{r}_{c}$, do not allow for every combination of charge and mass. For black holes in Minkowski space, this leads to the extremality bound $Q\leq M$. The constraint comes from the fact that the inner horizon cannot cross the outer horizon, i.e. $\hat{r}_{-}\ngeq\hat{r}_{+}$, preventing the development of a naked singularity. For the RN-dS black hole, the same reasoning goes true, although the extremality bound $Q\leq M$ is slightly altered. Additionally, the outer horizon of the black hole is not allowed to be bigger than the cosmological horizon. As the three radii are related through eq. (13), these constraints can be summarised into one inequality, which in terms of the charge and mass takes the following form:
@@
Figure 1: Only black holes with charge and mass under the curve are allowed. The top blue line corresponds to extremal black holes where inner and outer black hole horizons meet, and is referred to as the ‘extremal branch’. The vertical blue line represents the black holes in the Nariai limit, where the outer horizon and cosmological horizon meet, which is referred to as the ‘Nariai branch’.The dashed line represents $Q=M$, and illustrates that the extremality bound for RN-dS black holes deviates from the one for RN black holes in Minkowski space.
In the Nariai limit, one lets the cosmological horizon and outer black hole horizon come close together such that they become indistinguishable for observers. However, they cannot become strictly equal, and therefore it is useful to expand the metric for an observer that feels no force in between those two horizons, located at $\hat{r}=\hat{r}_{g}$ with $f^{\prime}(\hat{r}_{g})=0$. The Nariai limit can then be understood by bringing the two horizons close to this static observer, although not touching. In principle, one can find the solution to the equation $f^{\prime}(\hat{r}_{g})=0$ in the form $\hat{r}_{g}=\hat{r}_{g}(\hat{r}_{+},\hat{r}_{c})$. However, the expression becomes rather lengthy and it is not necessary for our discussion, so we do not write it here. So in what follows, we want to consider the limit $\hat{r}_{+}\rightarrow\hat{r}_{g}\leftarrow\hat{r}_{c}$. It proves to be useful to perform a coordinate transformation as follows:
@@
which have the same solutions as in eq. (24). Furthermore, the Nariai background is also a solution of “Einstein-Born-Infeld” theory Cai:2004eh , where one replaces the Maxwell action of electromagnetism with the Born-Infeld action, with a constant electric field.
As explained in the introduction, the goal of this paper is to embed the Nariai black hole into the dark bubble. Let us therefore end this section with some expectations on how the 5d embedding should look like. In the dark bubble model, any 4d energy density on the brane bends it outwards. The first example was discussed in Banerjee:2019fzz , where point particles in 4d correspond to strings that pull on the brane in the extra dimension. Additionally, in ref. Basile:2023tvh it was shown that in the presence of an electromagnetic field on the brane, there also exists a $B$-field that pulls the brane outwards.
@@ 4 The 5d black string
In particular, we start from five-dimensional Einstein gravity coupled to a three-form field strength $H$ with following action:
@@
where $\Lambda_{5}=-4k^{2}$ is the 5d cosmological constant with $k$ the inverse AdS5 length, and $H$ a three-form field strength with two-form gauge potential $B$ [^3] In a string theory context, the cosmological constant arises from moduli stabilisation. Notice also that we have omitted a dilatonic coupling to $H$. We justify this in Appendix A. For the metric, we take a modified black string Ansatz that makes the dS2 structure of the Nariai black hole that we are aiming for more apparent:
@@
When $k_{2}$ is turned on, the Einstein equations receive additional terms and hence the solutions change. First, eq. (36) informs us that $f=\tilde{f}$ and we can take $c^{\prime}=1$ as it can be absorbed by redefinition of the time coordinate $t$. This implies that we need to take $c_{t}=c_{r}$, and such 5d solutions turn out to be extremal at zeroth order in $k_{2}$. We then solve the system of differential equations perturbatively when $k_{2}^{2}/k^{2}$ becomes small. We can write the expansions
@@
When plugging these into the Einstein equations, we find the following Fefferman-Graham expressions of the first order corrections:
@@ 5.1 Solving the junction conditions
At this stage, we are ready to glue two different black string spacetimes together. They are joined together by sharing a boundary, which corresponds to a (dynamical) shell. In our setting, the shell is provided by a three-brane that is embedded at a fixed location in the $u$-direction, i.e. at $u=a$ with $a$ constant. The prescription for gluing spacetimes was originally developed by Israel in the form of junction conditions Israel:1966rt . The first condition requires the induced metric on the shell (three-brane) to be continuous:
@@
Now we have everything to solve the junction conditions, for two spacetimes joint together by a three-brane at $u=a$. The first junction condition requires us to rescale the time and radial coordinates as follows, defining new coordinates $t$ and $r$:
@@
$+\left(\frac{(1-12k_{-}^{2}q_{-}^{2})\log(k_{-}a)}{12k_{-}^{3}}-\frac{(1-12k_{+}^{2}q_{+}^{2})\log(k_{+}a)}{12k_{+}^{3}}\right)\frac{1}{a^{4}}$ (60)
@@ 5.2 Matching with 4d Nariai
In these expressions it is useful to identify the 4d Planck scale using the standard dark bubble relation
@@
Per usual, the Nariai system is solved for $a$ and $k_{2}$ in terms of $Q$ and $k_{4}$, but here we would like to solve for $a$ and $\sigma$ (or $k_{4}$) instead. We find that
$a^{2}=\frac{1-\sqrt{1-8k_{2}^{2}Q^{2}}}{2k_{2}^{2}},\qquad k_{4}^{2}=\frac{1+4k_{2}^{2}Q^{2}+\sqrt{1-8k_{2}^{2}Q^{2}}}{24Q^{2}}$ (77)
and hence
$\ell_{\text{Pl},5}^{3}\sigma=3(k_{-}-k_{+})-\frac{3}{2}\left(\frac{1}{k_{+}}-\frac{1}{k_{-}}\right)\frac{1+4k_{2}^{2}Q^{2}+\sqrt{1-8k_{2}^{2}Q^{2}}}{24Q^{2}}$ (78)
We need to check as a consistency requirement that $(k_{2}a)^{2}\ll 1$ and $\sigma>0$, and preferably that $\ell_{\text{Pl},5}^{3}\sigma-3(k_{-}-k_{+})\ll 1$. This is accounted for when $k_{2}^{2}Q^{2}\ll 1$, which can be arranged for with an appropriate choice of $c_{r+}$. We want to stress again that, although $q_{+}$ and $c_{r+}$ are not completely independent in the ordinary black string setup, we can take them independent here as we do not need to meet the requirements that the solution on the outside extrapolates to a solution for small values of $a$ without naked singularities.
@@ 6 Relating fine structure constant to the string coupling
There now exists a very non-trivial way to test the dark bubble model of the Nariai black hole. In ref. Banerjee:2019fzz , it was shown that a string pulling on the brane corresponds to a particle of mass $\tau_{\text{ps}}L$, where $\tau_{\text{ps}}$ is the effective tension of the string and $L=1/k$, the AdS radius. The effective tension can be obtained by integrating the energy density over the $S^{2}$. The difference in tension (including the energy density of the electric flux) compared to the critical tension contributes to the energy density as the cosmological constant.
@@
To test this condition, let us first compare our charge $Q$ with its usual definition in Maxwell theory in terms of the fine structure constant and fundamental charge. We make use of the assumptions at the end of the previous section and find using eq. (67)
@@ 7 Discussion
Next, one could wonder about Gregory-Laflamme instabilities Gregory:1993vy for the setup we consider. According to the Gubser-Mitra conjecture Gubser:2000ec ; Gubser:2000mm , a classical Gregory-Laflamme instability should correspond to a thermodynamic instability, i.e. when the specific heat becomes negative. In Bernamonti:2007bu , a critical value of the 5d charge was identified beyond which the specific heat is always positive. Below this value, three different branches open up of which two are stable. It turns out that the extremal black string solution we consider on the outside, i.e. with $c_{r}=c_{t}$, has zero Hawking temperature and belongs to a stable branch. For the inside, where we are also interested in extremal background with vanishing 5d charge, it was shown in ref. Brihaye:2007ju that the classical instability does arise for this regular background. A possible solution to this problem is to consider a non-vanishing 5d charge on the inside as well. Notice that all these observations are only studied for the case $k_{2}=0$. It would be interesting for the future to see how the situation is affected for non-vanishing $k_{2}$. Additionally, one should remember that these instabilities have been investigated for static backgrounds. In the full dark bubble setup that we have in mind, the black hole is more than the cylinder, as depicted in figure 2. The black hole is riding a brane that is expanding. It would be interesting to see if there is a connection between this and classical instability.
Next, we managed to relate the electromagnetic fine structure constant to the string coupling, in agreement with the calculation for an extremal black hole of unit charge, as in Danielsson:2023alz . Remarkably, the exact same relation is obtained even though our calculation for the cosmological Nariai black hole does not make use of the microscopic relations between, e.g., the string scale and the higher dimensional scale $L$. This suggests an internal consistency of the dark bubble that is replicated by string theory.

### v2

@@ 1 Introduction
According to the dark bubble model, the universe is riding an expanding bubble in AdS5, with the radial direction along the throat Banerjee:2018qey . The bubble is formed through a phase transition, where a bubble of true vacuum nucleates. Its wall consists of a brane with a tension slightly less than the critical value, with the difference setting the value of the positive cosmological constant. The model has been studied and developed in a series of papers, see refs. Banerjee:2018qey ; Banerjee:2019fzz ; Banerjee:2020wix ; Banerjee:2020wov ; Banerjee:2021qei ; Danielsson:2022fhd ; Danielsson:2022lsl ; Danielsson:2023alz ; Basile:2023tvh ; Banerjee:2023uto , with the earlier results reviewed in refs. Banerjee:2021yrb ; Banerjee:2022ree and related studies in Berglund:2022qsb ; Bandos:2023yyo ; Koga:2019yzj ; Koga:2020jok ; Koga:2022opd ; Basile:2020mpt ; Basile:2021vxh . In Danielsson:2022lsl ; Danielsson:2023alz the dark bubble model was embedded into a specific string theory setting, where the unique hierarchy of scales of the model was explored. In standard compactifications and RS-brane worlds, the 5d Planck length is always larger than the 4d one, provided that the size of the extra dimension, and the AdS length scale, are larger than the higher dimensional Planck length. The dark bubble, which has a brane with an inside and an outside, rather than the two insides of the RS-brane, is fundamentally different. What matters is the difference in AdS scale between the inside and the outside, which leads to a smaller 5d Planck length than the 4d one. This leads to several unique phenomenological features - including no need for scale separation in the higher dimensions.
@@
In this paper, we consider another case that further tests the dark bubble model: a (charged) black hole. We will restrict ourselves to the simplest setup, a charged Nariai black hole as large as the cosmological horizon [^1] While the Nariai solution is not relevant for a direct realistic application in cosmology, its embedding serves as an important check on the theoretical consistency of the dark bubble model. Nariai:1999 ; Romans:1991nq . Such black holes have been studied in the context of the Swampland paradigm Montero:2019ekk , where it has been argued that their existence leads to new constraints on particle masses beyond those of the weak gravity conjecture. In this paper, we show that the dark bubble scenario can successfully incorporate a Nariai black hole. This specific construction serves as a stepping stone for realising more general black holes on the dark bubble. The analysis provides yet another explicit example of how the dark bubble reproduces 4d Einstein gravity without localising it on the brane, a characteristic of the RS model. Instead, the bulk plays an important role in the construction, and the failure to localise yields well-controlled corrections that become important only at stringy densities, Basile:2023tvh .
@@ 2 The dark bubble scenario
and $e^{\alpha}_{\;b}$ are the tangent vectors to the brane for every $b$. The “--” and “$+$” correspond to the contributions from the in- and outside spacetimes respectively. The 4d Newton’s constant is related to the 5d one through
@@
An important property, peculiar to the dark bubble, is that $S_{ab}$ seemingly contributes with the wrong sign to the 4d energy density. This is the minus sign in front of $S_{ab}$ in (2). In the case of a pure brane, we simply have $S_{ab}=-\sigma h_{ab}$, which is why a larger value of the brane density $\sigma$ reduces the effective cosmological constant given by $\Lambda_{4}=\sigma_{c}-\sigma$ [^3] Note that we are using $+$-$+$- + + + signature so that $h_{tt}<0$ has $S_{tt}>0$ (and $S^{t}_{t}<0$) for a positive energy density. For the dark bubble to nucleate, its tension needs to be smaller than the critical value $\sigma_{c}$, guaranteeing a positive cosmological constant.
To get the full energy-momentum tensor $T_{ab}$ in the presence of matter on the brane, one must take into account its backreaction on the bulk, captured by $\mathcal{J}_{ab}$. This gives rise to an effective total energy density in 4d that is equal to the physical one. It is this phenomenon that has been exhibited in earlier work mentioned in the introduction. The choice of the bulk solution is part of setting up the appropriate initial conditions for the physical situation of interest. It is a non-trivial result of ref. Basile:2023tvh that this can be done for electromagnetic waves such that the subsequent time evolution is that of general relativity coupled to Maxwell theory. A crucial aspect is how the electromagnetic field on the brane sources a string $B$-field in the bulk. The backreaction of this field on the bulk metric is what allows the tensor $\mathcal{J}_{ab}$ to give rise to the desired energy-momentum tensor $T_{ab}$.
@@ 3 Reissner-Nordström-de Sitter black holes and the Nariai limit
In this section, we briefly review Reissner-Nordström-de Sitter (RN-dS) black holes and the Nariai limit in four dimensions. These have been receiving increased attention recently and led to the Festina Lente bound in Montero:2019ekk . These black holes are solutions of Einstein-Maxwell theory given by the familiar action
@@
The constraint (13) and the fact that $\hat{r}_{-}\leq\hat{r}_{+}\leq\hat{r}_{c}$, do not allow for every combination of charge and mass. For black holes in Minkowski space, this leads to the extremality bound $Q\leq M$. The constraint comes from the fact that the inner horizon cannot cross the outer horizon, i.e. $\hat{r}_{-}\ngeq\hat{r}_{+}$, preventing the development of a naked singularity. For the RN-dS black hole, the same reasoning goes true, although the extremality bound $Q\leq M$ is slightly altered. Additionally, the outer horizon of the black hole is not allowed to be bigger than the cosmological horizon. As the three radii are related through eq. (13), these constraints can be summarised into one inequality, which in terms of the charge and mass takes the following form Romans:1991nq :
@@
Figure 1: Only black holes with charge and mass under the curve are allowed. The top blue line corresponds to extremal black holes where inner and outer black hole horizons meet, and is referred to as the ‘extremal branch’. The vertical blue line represents the black holes in the Nariai limit, where the outer horizon and cosmological horizon meet, which is referred to as the ‘Nariai branch’.The dashed line represents $Q=M$ and illustrates that the extremality bound for RN-dS black holes deviates from the one for RN black holes in Minkowski space.
In the Nariai limit, one lets the cosmological horizon and outer black hole horizon come close together such that they become indistinguishable for observers. However, they cannot become strictly equal, and therefore it is useful to expand the metric for an observer that feels no force in between those two horizons, located at $\hat{r}=\hat{r}_{g}$ with $f^{\prime}(\hat{r}_{g})=0$. The Nariai limit can then be understood by bringing the two horizons close to this static observer, although not touching. In principle, one can find the solution to the equation $f^{\prime}(\hat{r}_{g})=0$ in the form $\hat{r}_{g}=\hat{r}_{g}(\hat{r}_{+},\hat{r}_{c})$. However, the expression becomes rather lengthy and it is not necessary for our discussion, so we do not write it here. So in what follows, we want to consider the limit $\hat{r}_{+}\rightarrow\hat{r}_{g}\leftarrow\hat{r}_{c}$, see also refs. Hawking:1995ap ; Bousso:1996au . It proves to be useful to perform a coordinate transformation as follows:
@@
which have the same solutions as in eq. (24). Furthermore, the Nariai background is also a solution of “Einstein-Born-Infeld” theory Cai:2004eh , where one replaces the Maxwell action of electromagnetism with the Born-Infeld action, with a constant electric field. Additionally, the role of dilatonic couplings has also been studied for the charged Nariai black hole setup Bousso:1996pn .
As explained in the introduction, the goal of this paper is to embed the Nariai black hole into the dark bubble. Let us therefore end this section with some expectations on how the 5d embedding should look. In the dark bubble model, any 4d energy density on the brane bends it outwards. The first example was discussed in Banerjee:2019fzz , where point particles in 4d correspond to strings that pull on the brane in the extra dimension. Additionally, in ref. Basile:2023tvh it was shown that in the presence of an electromagnetic field on the brane, there also exists a $B$-field that pulls the brane outwards.
@@ 4 The 5d black string
In particular, we start from five-dimensional Einstein gravity coupled to a three-form field strength $H$ with the following action:
@@
where $\Lambda_{5}=-4k^{2}$ is the 5d cosmological constant with $k$ the inverse AdS5 length, and $H$ a three-form field strength with two-form gauge potential $B$ [^4] In a string theory context, the cosmological constant arises from moduli stabilisation. Notice also that we have omitted a dilatonic coupling to $H$. We have justified this in Appendix A, where it is shown that a non-trivial dilaton can be neglected in the approximation scheme used in this paper. For the metric, we take a modified black string Ansatz that makes the dS2 structure of the Nariai black hole that we are aiming for more apparent:
@@
When $k_{2}$ is turned on, the Einstein equations receive additional terms and hence the solutions change. First, eq. (36) informs us that $f=c^{\prime}\tilde{f}$ and we can take $c^{\prime}=1$ as it can be absorbed by redefinition of the time coordinate $t$. This implies that we need to take $c_{t}=c_{r}$, and such 5d solutions turn out to be extremal at zeroth order in $k_{2}$. We then solve the system of differential equations perturbatively when $k_{2}^{2}/k^{2}$ becomes small. We can write the expansions
@@
When plugging these into the Einstein equations, we find the following Fefferman-Graham expressions of the first-order corrections:
@@ 5.1 Solving the junction conditions
At this stage, we are ready to glue two different black string spacetimes together. They are joined together by sharing a boundary, which corresponds to a (dynamical) shell. In our setting, the shell is provided by a three-brane that is embedded at a fixed location in the $u$-direction, i.e. at $u=a$ with $a$ constant. The prescription for glueing spacetimes was originally developed by Israel in the form of junction conditions Israel:1966rt . The first condition requires the induced metric on the shell (three-brane) to be continuous:
@@
Now we have everything to solve the junction conditions, for two spacetimes joined together by a three-brane at $u=a$. The first junction condition requires us to rescale the time and radial coordinates as follows, defining new coordinates $t$ and $r$:
@@
$+\left(\frac{(1-12k_{-}^{2}q_{-}^{2})\log(k_{-}a)}{1k_{-}^{3}}-\frac{(1-12k_{+}^{2}q_{+}^{2})\log(k_{+}a)}{12k_{+}^{3}}\right)\frac{1}{a^{4}}$ (60)
@@ 5.2 Matching with 4d Nariai
In these expressions, it is useful to identify the 4d Planck scale using the standard dark bubble relation
@@
Per usual, the Nariai system is solved for $a$ and $k_{2}$ in terms of $Q$ and $k_{4}$. We find that
$k_{2}^{2}a^{2}=\sqrt{1-12k_{4}^{2}Q^{2}}\qquad k_{4}^{2}=\frac{1+k_{2}^{2}a^{2}}{6a^{2}}$ (77)
and for the tension, we have that
$\ell_{\text{Pl},5}^{3}\sigma=3(k_{-}-k_{+})-\frac{1}{2}\left(\frac{1}{k_{+}}-\frac{1}{k_{-}}\right)3k_{4}^{2}=3(k_{-}-k_{+})\left(1-\frac{k_{4}^{2}}{2k_{-}k_{+}}\right)$ (78)
We need to check as a consistency requirement that $(k_{2}a)^{2}\ll 1$ and $\sigma>0$, and preferably that $\ell_{\text{Pl},5}^{3}\sigma-3(k_{-}-k_{+})\ll 1$. The first requirement stems from eq. (55) and is accounted for whenever $1-12(k_{4}Q)^{2}\ll 1$. This means that the Nariai black holes we are describing, are near-extremal, and hence the chargeless-limit $k_{4}Q\to 0$ cannot be taken. Additionally, one sees that for the near-extremal limit, $k_{4}^{2}\approx 1/(6a^{2})$. Hence $k_{4}^{2}/(2k_{+}k_{-})\approx 1/(12(k_{+}a)(k_{-}a))\ll 1$ and the tension is indeed positive and close to the critical tension. The near-extremal Nariai limit can be obtained by an appropriate choice of $c_{r+}$. We want to stress again that, although $q_{+}$ and $c_{r+}$ are not completely independent in the ordinary black string setup, we can take them independent here as we do not need to meet the requirements that the solution on the outside extrapolates to a solution for small values of $a$ without naked singularities.
@@ 6 Relating fine structure constant to the string coupling
There exists a very non-trivial way to test the dark bubble model of the Nariai black hole. In ref. Banerjee:2019fzz , it was shown that a string pulling on the brane corresponds to a particle of mass $\tau_{\text{ps}}L$, where $\tau_{\text{ps}}$ is the effective tension of the string and $L=1/k$, the AdS radius. The effective tension can be obtained by integrating the energy density over the $S^{2}$. The difference in tension (including the energy density of the electric flux) compared to the critical tension contributes to the energy density as the cosmological constant.
@@
To test this condition, let us first compare our charge $Q$ with its usual definition in Maxwell’s theory in terms of the fine structure constant and fundamental charge. We make use of the assumptions at the end of the previous section and find using eq. (67)
@@ 7 Discussion
Next, one could wonder about Gregory-Laflamme instabilities Gregory:1993vy for the setup we consider. According to the correlated stability conjecture Gubser:2000ec ; Gubser:2000mm , a classical Gregory-Laflamme instability should correspond to a thermodynamic instability, i.e. when the specific heat becomes negative. This has been verified for many examples, although counterexamples have been found as well where the black strings have scalar hair, see e.g. refs. Friess:2005zp ; Buchel:2010wk ; Buchel:2011ra . In Bernamonti:2007bu , where no scalar hair is present and hence the conjecture is expected to hold, a critical value of the 5d charge was identified beyond which the specific heat is always positive. Below this value, three different branches open up of which two are stable. It turns out that the extremal black string solution we consider on the outside, i.e. with $c_{r}=c_{t}$, has zero Hawking temperature and belongs to a stable branch. For the inside, where we are also interested in extremal background with vanishing 5d charge, it was shown in ref. Brihaye:2007ju that the classical instability does arise for this regular background. A possible solution to this problem is to consider a non-vanishing 5d charge on the inside as well. Notice that all these observations are only studied for the case $k_{2}=0$. It would be interesting to see how the situation is affected by non-vanishing $k_{2}$ in the future. Additionally, one should remember that these instabilities have been investigated for static backgrounds. In the full dark bubble setup that we have in mind, the black hole is more than the cylinder, as depicted in figure 2. The black hole is riding a brane that is expanding. It would be interesting to see if there is a connection between this and classical instability.
Next, we managed to relate the electromagnetic fine structure constant to the string coupling, in agreement with the calculation for an extremal black hole of unit charge, as in Danielsson:2023alz . Remarkably, the same relation is obtained even though our calculation for the cosmological Nariai black hole does not make use of the microscopic relations between, e.g., the string scale and the higher dimensional scale $L$. This suggests an internal consistency of the dark bubble that is replicated by string theory.